# A Hybrid Quantum-assisted Column Generation Algorithm for the Fleet Conversion Problem

### 11 Jan 2024 v2

@@
Abstract.
The problem of Fleet Conversion aims to reduce the carbon emissions and cost of operating a fleet of vehicles for a given set of tours. It can be modelled as a column generation scheme with the Maximum Weighted Independent Set(MWIS) problem as the slave. Quantum variational algorithms have gained significant interest in the past several years. Recently, a method to represent Quadratic Unconstrained Binary Optimization(QUBO) problems using logarithmically fewer qubits was proposed. Here we use this method to solve the MWIS Slaves and demonstrate how quantum and classical solvers can be used together to approach an industrial-sized use-case (up to 64 tours).
quantum variational algorithms; optimization; column generation
copyright: acmcopyrightjournalyear: 2023doi: XXXXXXX.XXXXXXXprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Transportationccs: Computer systems organization Quantum computing
1. Introduction
@@
Quantum computing (Nielsen and Chuang, 2011; Preskill, 2021, 2018) is a potentially disruptive field that could have applications in several domains including financial modelling (Orús et al., 2019; Herman et al., 2022), cryptography (Scarani et al., 2009; Mehic et al., 2020), chemistry (Martin et al., 2022; Haidar et al., 2023) and optimization. Within the scope of optimization applications, there has been a growing interest in quantum variational algorithms (Bittel and Kliesch, 2021; Cerezo et al., 2021; Callison and Chancellor, 2022; Peruzzo et al., 2014; Nakanishi et al., 2020; Moll et al., 2018; Lubasch et al., 2020; Stokes et al., 2020). Among them, the Quantum Approximate Optimization Algorithm (QAOA) has been heavily researched (Farhi et al., 2014; Fuchs et al., 2021; Herrman et al., 2021; Larkin et al., 2022; Zhou et al., 2020; Hadfield et al., 2019). A well known issue with QAOA is that it does not scale well with problem size limiting its applications to toy problems . Recently, an algorithm to treat Quadratic Unconstrained Binary Optimization (QUBO) (Glover et al., 2019; Date et al., 2021; Calude et al., 2017; Papalitsas et al., 2019) problem using logarithmically fewer qubits has been demonstrated(Rančić, 2023; Chatterjee et al., 2023). In this paper, we use column generation (Desaulniers et al., 2006; Barnhart et al., 1998; Wilhelm, 2001; Demiriz et al., 2002; Mehrotra and Trick, 1996) to describe our problem as a master problem and several sub-problems henceforth referred to as slaves. The slave problem in our case is the Maximum Weighted Independent Set (MWIS) problem which can be represented as a QUBO problem. We propose an algorithm that handles the master problem using a commercial linear program solver Gurobi and the slave problems using a quantum solver based on (Chatterjee et al., 2023). In our experiments, we solve instances up to a size of 64 tours using only 7 qubits to represent the MWIS Slaves. This shows that the method is compatible with the quantum computers of the NISQ era.
@@
2. Problem Statement and Methods
@@ 2.1. Statement
Each tour $k\in K$ is described by the tuple $(t_{d}^{k},t_{a}^{k},l_{d}^{k},l_{a}^{k},A^{k})$, corresponding respectively to the departure time, arrival time, departure location and arrival location, and a set of authorized vehicle models of the tour. The time to travel from location $i$ to location $j$ ($TT_{ij}$) can be computed using the distance matrix of the locations. This matrix is used to derive the time needed to relocate any physical vehicle from the arrival location of a tour to the departure location of another tour in case these tours are meant to be assigned to the same color.
@@
(1) $x^{k}_{c}=\left\{\begin{matrix}1&\mbox{if color $c$ is assigned to tour $k$}\\ 0&\mbox{otherwise,}\end{matrix}\right.$
@@
(2) $y_{c}=\left\{\begin{matrix}1&\mbox{if color $c$ is purchased}\\ 0&\mbox{otherwise.}\end{matrix}\right.$
@@
(3) $\min$ $\sum_{c\in C}\gamma_{c}y_{c}+\sum_{k\in K,c\in C}\Gamma^{k}_{c}x_{c}^{k}$
(4) s.t. $\sum_{c}x^{k}_{c}\geq 1$ $(\forall k\in K)$
(5) $x^{k}_{c}\leq y_{c}$ $(\forall c\in C)(\forall k\in K)$
(6) $x^{k}_{c}+x^{k^{\prime}}_{c}\leq 1$ ($(\forall\{k,k^{\prime}\}\in\mathcal{I})(\forall c\in C)$
(7) $x^{k}_{c}=0$ $(\forall k\in K)(\forall c\in C)(v(c)\notin A^{k})$
(8) $y_{c},x_{c}^{k}\in\{0,1\}.$
@@
2.2. Formal description as a Graph Problem
@@
Figure 1. Incompatibility Graph generated from the time windows of tours. The lines represent the time windows of 7 tours. Colors blue, red and green are used to assign vehicles to tours taking into account their incompatibility.
@@
(9) $\Lambda=\{(a_{I},c_{I})|I\mbox{ independent set of }G|\mbox{ s.t. }\forall k\in K,a^{k}_{I}=1\Rightarrow v(c_{I})\in A^{k}\}.$
@@
(10) $\min$ $\sum_{I\in\Lambda}\Gamma_{I}x_{I}$
(11) s.t. $\sum_{I\in\Lambda}a_{I}^{k}x_{I}\geq 1$ $(\forall k\in K)$
(12) $x_{I}=\begin{cases}1,\text{ if independent set }I\text{ chosen }\\ 0,\text{ otherwise}\end{cases}$
@@
With the above reformulation, the problem looks considerably smaller. We have indeed only one constraint per tour. However, we need to generate the set $\Lambda$ in order to solve this problem. The set $\Lambda$ contains all the feasible independent sets of the graph, that is, independent set of the graph coupled with colors that are compatible with all nodes therein. Remember that if $I$ is a non-empty independent set of $G$ and $c\neq c^{\prime}$ are two different colors, then $(I,c)$ and $(I,c^{\prime})$ are two distinct elements of $\Lambda$ as they describe two different conversion solutions. We thus have a very large number of variables $x_{I}$. Nevertheless, it is clear that only a small number of those variables should be non-zero at optimum. Indeed, at worst (in terms of number of independent sets activated), no sharing of color is possible and each tour has a dedicated vehicle, which corresponds to $|K|$ variables $x_{I}$ activated. This is why an interesting approach here is column generation, where independent sets are generated dynamically while the solution converges to an optimum.
2.3. The Column Generation Algorithm
In this paragraph, we design the column generation procedure to solve our problem. First, we derive an extension of the problem that makes it always feasible. To do so, we permit the algorithm to reject tours from the solution. Second, we relax all integrity constraints and actually solve the LP-relaxation of our problem. Once an LP-optimal solution is found by the column generation, one can summon any type of rounding algorithm to build (if needed) an integer-feasible solution from the relaxed solution. For instance, see (Bertsimas et al., 1999). We focus this work on the problem of finding the optimal LP-feasible solution for our problem.
@@ 2.3. The Column Generation Algorithm
(13) $r_{k}=\left\{\begin{matrix}1&\mbox{if }k\mbox{ is rejected}\\ 0&\mbox{otherwise.}\end{matrix}\right.$
@@
(MP) $\min$ $\sum_{I\in\Lambda}\Gamma_{I}x_{I}+R\sum_{k\in K}r_{k}$
(MP-1) s.t. $\sum_{I\in\Lambda}a_{I}^{k}x_{I}+r_{k}\geq 1$ $(\forall k\in K)$
(MP-2) $x_{I},r_{k}\geq 0$ $(\forall I\in\Lambda)(\forall k\in K)$
When the binary constraints are enforced on $x_{I}$ and $r_{k}$, and if $R$ is sufficiently large, it is clear that Problems (MP) and (10)–(11) are equivalent. The dual program (Balinski and Tucker, 1969) of (MP) reads :
(D) $\max$ $\sum_{k\in K}\mu_{k}$
(D-1) s.t. $\sum_{k\in K}a^{k}_{I}\mu_{k}\leq\Gamma_{I}$ $(\forall I\in\Lambda)$
(D-2) $\mu_{k}\geq 0$ $(\forall k\in K)$
@@
(RMP) $\min$ $\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x_{I}+R\sum_{k\in K}r_{k}$
(RMP-1) s.t. $\sum_{I\in\Lambda^{\prime}}a_{I}^{k}x_{I}+r_{k}\geq 1$ $(\forall k\in K)$
(RMP-2) $x_{I},r_{k}\geq 0$ $(\forall I\in\Lambda^{\prime})(\forall k\in K)$
(RD) $\max$ $\sum_{k\in K}\mu_{k}$
(RD-1) s.t. $\sum_{k\in K}a^{k}_{I}\mu_{k}\leq\Gamma_{I}$ $(\forall I\in\Lambda^{\prime})$
(RD-2) $\mu_{k}\geq 0$ $(\forall k\in K)$
Let $F$ (resp. $F^{\prime}$) denote the feasible set of (MP) (resp. (RMP)) and $G$ (resp. $G^{\prime}$) be the feasible set of (D) (resp. (RD)). It is clear that $F^{\prime}\subset F$ and $G\subset G^{\prime}$. Furthermore, for any $\Lambda^{\prime}\subset\Lambda$, (RMP) is linear, feasible and bounded. Therefore, strong-duality applies (Balinski and Tucker, 1969)(Gale et al., 1951). This means that (RMP) and (RD) are both feasible and a primal-dual pair of solutions $(x^{\prime},\mu^{\prime})$ exists, where $x^{\prime}$ solves (RMP), $\mu^{\prime}$ solves (RD). Furthermore, it means we have the equality:
(14) $\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x^{\prime}_{I}+R\sum_{k\in K}r^{\prime}_{k}=\sum_{k\in K}\mu^{\prime}_{k}.$
@@
Fact 2 ().
Let $\Lambda^{\prime}\subset\Lambda$. Let $(x^{*},\mu^{*})$ and $(x^{\prime},\mu^{\prime})$ be primal-dual optimal couples for (MP)–(D) and (RMP)–(RD) respectively. By definition:
@@
  • $\mu^{*}\in G\subset G^{\prime}$
  • $\mu^{\prime}\in G^{\prime}$
Suppose that $\mu^{\prime}\in G$.
Then, $x^{\prime}\in F$ and $x^{\prime}$ is optimal for (MP).
Indeed, if $\mu^{\prime}\in G$, by definition of $\mu^{*}$, we know that $\sum_{k\in K}\mu^{\prime}_{k}\leq\sum_{k\in K}\mu^{*}_{k}$. On the other hand, as $G\subset G^{\prime}$, by definition of $\mu^{\prime}$, we have $\sum_{k\in K}\mu^{\prime}_{k}\geq\sum_{k\in K}\mu^{*}_{k}$. Thus, we have equality. By strong duality, this means that
(15) $\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x^{\prime}_{I}+R\sum_{k\in K}r^{\prime}_{k}$ $=\sum_{k\in K}\mu^{\prime}_{k}$
(16) $=\sum_{k\in K}\mu^{*}_{k}$
(17) $=\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x^{*}_{I}+R\sum_{k\in K}r^{*}_{k}$
Therefore, if $\mu^{\prime}$ is feasible in (D), then $x^{\prime}$ is an optimal solution of (MP).
@@
(18) $\sum\limits_{k}a^{k}_{I}\mu^{\prime}_{k}\leq\Gamma_{I}\text{ }\forall\text{ }I\in\Lambda$
Therefore, the existence of violated constraints (18) means that the current $x^{\prime}$ is not the optimum and that new columns (allocations) can be added to improve the solution. We can therefore try to find an independent set that minimizes the reduced cost $\Gamma_{I}-\sum\limits_{k}y^{k}_{I}\mu^{\prime}_{k}$. If this reduced cost is negative then the independent set $I$ used to obtain this negative reduced cost violates (18) and can therefore be added to the set of independent sets in RMP $\Lambda^{\prime}$.
@@
(20) $\begin{split}\sigma&=\Gamma_{I}-\sum\limits_{k}a^{k}_{I}\mu^{\prime}_{k}\\ &=\sum\limits_{k}a_{I}^{k}\Gamma_{c}^{k}-\sum\limits_{k}a^{k}_{I}\mu^{\prime}_{k}\\ &=\sum\limits_{k}a_{I}^{k}(\Gamma_{c}^{k}-\mu^{\prime}_{k})\end{split}$
@@
(21) $\max\sigma=\sum\limits_{k}a_{I}^{k}(\mu^{\prime}_{k}-\Gamma_{c}^{k})$
By definition, $(a_{I}^{k})_{k\in K}$ is a vector denoting an independent set, and $\mu^{\prime}_{k}-\Gamma_{c}^{k}$ can be seen as numerical weights for every color $c$. This is therefore a Maximum Weighted Independent Set problem. For the rest of the paper, this problem will be our slave problem (SP). The MWIS problem can be defined as follows.
(SP) $\max\sigma=$ $\sum_{k}y_{k}(\mu^{\prime}_{k}-\Gamma_{c}^{k})$
(SP-1) s.t. $y_{k}+y_{j}\leq 1$ $(\forall(k,j)\in E)$
(SP-2) $y_{k}\in\{0,1\}$ $(\forall k\in K)$
A solution of Problem (SP) defines an allocation $(a_{I},c)$ where $I\coloneqq\{k\in K|y_{k}=1\}$ and $a_{I}^{k}=y_{k}$. Note that there is one slave problem per color, and, as in our problem, colors from the same model are equivalent, there is actually one problem per vehicle model. Therefore, each allocation produced by a slave represents a single physical vehicle (color) along with all the tours assigned to it. Given the above problem definitions, we can define the algorithm to solve RMP as described in Algorithm 1.

### 12 Mar 2024 v3

@@
Abstract
The problem of Fleet Conversion aims to reduce the carbon emissions and cost of operating a fleet of vehicles for a given set of tours. It can be modelled as a column generation scheme with the Maximum Weighted Independent Set(MWIS) problem as the sub-problem or worker problem. Quantum variational algorithms have gained significant interest in the past several years. Recently, a method to represent Quadratic Unconstrained Binary Optimization(QUBO) problems using logarithmically fewer qubits was proposed. Here we use this method to solve the MWIS Workers and demonstrate how quantum and classical solvers can be used together to approach an industrial-sized use-case (up to 64 tours).
keywords:
quantum variational algorithms, optimization, column generation
1 Introduction
@@
Quantum computing [1, 2, 3] is a potentially disruptive field that could have applications in several domains including financial modelling [4, 5], cryptography [6, 7], chemistry [8, 9] and optimization. Within the scope of optimization applications, there has been a growing interest in quantum variational algorithms [10, 11, 12, 13, 14]. Among them, the Quantum Approximate Optimization Algorithm (QAOA) has been heavily researched [15, 16, 17]. A well known issue with QAOA is that it does not scale well with problem size limiting its applications to toy problems. Recently, an algorithm to treat Quadratic Unconstrained Binary Optimization (QUBO) [18, 19] problem using logarithmically fewer qubits has been demonstrated[20, 21]. In this paper, we use column generation [22, 23] to describe our problem as a coordinator problem and several sub-problems henceforth referred to as workers. The worker problem in our case is the Maximum Weighted Independent Set (MWIS) problem which can be represented as a QUBO problem. We propose an algorithm that handles the coordinator problem using a commercial linear program solver Gurobi and the worker problems using a quantum solver based on [21]. In our experiments, we solve instances up to a size of 64 tours using only 7 qubits to represent the MWIS workers. This shows that the method is compatible with the quantum computers of the Noisy Intermediate Scale Quantum (NISQ) era, where the quantum computers are still relatively small and are not free of noise.
@@
2 Problem Statement and Methods
@@ 2.1 Statement
Each tour $k\in K$ is described by the tuple $(t_{d}^{k},t_{a}^{k},l_{d}^{k},l_{a}^{k},A^{k})$, corresponding respectively to the departure time, arrival time, departure location and arrival location, and a set of authorized vehicle models of the tour. The time to travel from location $i$ to location $j$ ($TT_{ij}$) can be computed using the distance matrix of the locations. This matrix is used to derive the compatibility time needed to relocate any physical vehicle from the arrival location of a tour to the departure location of another tour in case these tours are meant to be assigned to the same color.
@@
$x^{k}_{c}=\left\{\begin{matrix}1&\mbox{if color $c$ is assigned to tour $k$}\\ 0&\mbox{otherwise,}\end{matrix}\right.$ (1)
@@
$y_{c}=\left\{\begin{matrix}1&\mbox{if color $c$ is purchased}\\ 0&\mbox{otherwise.}\end{matrix}\right.$ (2)
@@
$\min$ $\sum_{c\in C}\gamma_{c}y_{c}+\sum_{k\in K,c\in C}\Gamma^{k}_{c}x_{c}^{k}$ (3)
s.t. $\sum_{c}x^{k}_{c}\geq 1$ $(\forall k\in K)$ (4)
$x^{k}_{c}\leq y_{c}$ $(\forall c\in C)(\forall k\in K)$ (5)
$x^{k}_{c}+x^{k^{\prime}}_{c}\leq 1$ ($(\forall\{k,k^{\prime}\}\in\mathcal{I})(\forall c\in C)$ (6)
$x^{k}_{c}=0$ $(\forall k\in K)(\forall c\in C)(v(c)\notin A^{k})$ (7)
$y_{c},x_{c}^{k}\in\{0,1\}.$ (8)
@@
2.2 Formal description as a Graph Problem
@@
Figure 1: Incompatibility Graph generated from the time windows of tours. The lines represent the time windows of 7 tours. Colors blue, red and green are used to assign vehicles to tours taking into account their incompatibility.
@@
$\Lambda=\{(a_{I},c_{I})|I\mbox{ independent set of }G|\mbox{ s.t. }\forall k\in K,a^{k}_{I}=1\Rightarrow v(c_{I})\in A^{k}\}.$ (9)
@@
$\min$ $\sum_{I\in\Lambda}\Gamma_{I}x_{I}$ (10)
s.t. $\sum_{I\in\Lambda}a_{I}^{k}x_{I}\geq 1$ $(\forall k\in K)$ (11)
$x_{I}=\begin{cases}1,\text{ if independent set }I\text{ chosen }\\ 0,\text{ otherwise}\end{cases}$ (12)
@@
With the above reformulation, the problem looks considerably smaller. We have indeed only one constraint per tour. However, we need to generate the set $\Lambda$ in order to solve this problem. The set $\Lambda$ contains all the feasible independent sets of the graph, that is, independent set of the graph coupled with colors that are compatible with all nodes therein. Remember that if $I$ is a non-empty independent set of $G$ and $c\neq c^{\prime}$ are two different colors, then $(I,c)$ and $(I,c^{\prime})$ are two distinct elements of $\Lambda$ as they describe two different conversion solutions. We thus have a very large number of variables $x_{I}$. Nevertheless, it is clear that only a small number of those variables should be non-zero at optimum. Indeed, at worst (in terms of number of independent sets activated), no sharing of color is possible and each tour has a dedicated vehicle, which corresponds to $|K|$ variables $x_{I}$ activated. This is why an interesting approach here is column generation [22], where independent sets are generated dynamically while the solution converges to an optimum.
2.3 The Column Generation Algorithm
In this paragraph, we design the column generation procedure to solve our problem. First, we derive an extension of the problem that makes it always feasible. To do so, we permit the algorithm to reject tours from the solution. Second, we relax all integrity constraints and actually solve the LP-relaxation of our problem. Once an LP-optimal solution is found by the column generation, one can summon any type of rounding algorithm to build (if needed) an integer-feasible solution from the relaxed solution. For instance, see [24]. We focus this work on the problem of finding the optimal LP-feasible solution for our problem.
@@ 2.3 The Column Generation Algorithm
$r_{k}=\left\{\begin{matrix}1&\mbox{if }k\mbox{ is rejected}\\ 0&\mbox{otherwise.}\end{matrix}\right.$ (13)
@@
$\min$ $\sum_{I\in\Lambda}\Gamma_{I}x_{I}+R\sum_{k\in K}r_{k}$ (CP)
s.t. $\sum_{I\in\Lambda}a_{I}^{k}x_{I}+r_{k}\geq 1$ $(\forall k\in K)$ (CP-1)
$x_{I},r_{k}\geq 0$ $(\forall I\in\Lambda)(\forall k\in K)$ (CP-2)
When the binary constraints are enforced on $x_{I}$ and $r_{k}$, and if $R$ is sufficiently large, it is clear that Problems (CP) and (10)–(11) are equivalent. The dual program [25] of (CP) reads :
$\max$ $\sum_{k\in K}\mu_{k}$ (D)
s.t. $\sum_{k\in K}a^{k}_{I}\mu_{k}\leq\Gamma_{I}$ $(\forall I\in\Lambda)$ (D-1)
$\mu_{k}\geq 0$ $(\forall k\in K)$ (D-2)
@@
$\min$ $\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x_{I}+R\sum_{k\in K}r_{k}$ (RCP)
s.t. $\sum_{I\in\Lambda^{\prime}}a_{I}^{k}x_{I}+r_{k}\geq 1$ $(\forall k\in K)$ (RCP-1)
$x_{I},r_{k}\geq 0$ $(\forall I\in\Lambda^{\prime})(\forall k\in K)$ (RCP-2)
$\max$ $\sum_{k\in K}\mu_{k}$ (RD)
s.t. $\sum_{k\in K}a^{k}_{I}\mu_{k}\leq\Gamma_{I}$ $(\forall I\in\Lambda^{\prime})$ (RD-1)
$\mu_{k}\geq 0$ $(\forall k\in K)$ (RD-2)
Let $F$ (resp. $F^{\prime}$) denote the feasible set of (CP) (resp. (RCP)) and $D$ (resp. $D^{\prime}$) be the feasible set of (D) (resp. (RD)). It is clear that $F^{\prime}\subset F$ and $D\subset D^{\prime}$. Furthermore, for any $\Lambda^{\prime}\subset\Lambda$, (RCP) is linear, feasible and bounded. Therefore, strong-duality applies [25][26]. This means that (RCP) and (RD) are both feasible and a primal-dual pair of solutions $(x^{\prime},\mu^{\prime})$ exists, where $x^{\prime}$ solves (RCP), $\mu^{\prime}$ solves (RD). Furthermore, it means we have the equality:
$\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x^{\prime}_{I}+R\sum_{k\in K}r^{\prime}_{k}=\sum_{k\in K}\mu^{\prime}_{k}.$ (14)
@@
Fact 2
Let $\Lambda^{\prime}\subset\Lambda$. Let $(x^{*},\mu^{*})$ and $(x^{\prime},\mu^{\prime})$ be primal-dual optimal couples for (CP)–(D) and (RCP)–(RD) respectively. By definition:
@@
  • $\mu^{*}\in D\subset D^{\prime}$
  • $\mu^{\prime}\in D^{\prime}$
Suppose that $\mu^{\prime}\in D$.
Then, $x^{\prime}\in F$ and $x^{\prime}$ is optimal for (CP).
Indeed, if $\mu^{\prime}\in D$, by definition of $\mu^{*}$, we know that $\sum_{k\in K}\mu^{\prime}_{k}\leq\sum_{k\in K}\mu^{*}_{k}$. On the other hand, as $D\subset D^{\prime}$, by definition of $\mu^{\prime}$, we have $\sum_{k\in K}\mu^{\prime}_{k}\geq\sum_{k\in K}\mu^{*}_{k}$. Thus, we have equality. By strong duality, this means that
$\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x^{\prime}_{I}+R\sum_{k\in K}r^{\prime}_{k}$ $=\sum_{k\in K}\mu^{\prime}_{k}$ (15)
$=\sum_{k\in K}\mu^{*}_{k}$ (16)
$=\sum_{I\in\Lambda^{\prime}}\Gamma_{I}x^{*}_{I}+R\sum_{k\in K}r^{*}_{k}$ (17)
Therefore, if $\mu^{\prime}$ is feasible in (D), then $x^{\prime}$ is an optimal solution of (CP).
@@
$\sum\limits_{k}a^{k}_{I}\mu^{\prime}_{k}\leq\Gamma_{I}\text{ }\forall\text{ }I\in\Lambda$ (18)
Therefore, the existence of violated constraints (18) means that the current $x^{\prime}$ is not the optimum and that new columns (allocations) can be added to improve the solution. We can therefore try to find an independent set that minimizes the reduced cost $\Gamma_{I}-\sum\limits_{k}y^{k}_{I}\mu^{\prime}_{k}$. If this reduced cost is negative then the independent set $I$ used to obtain this negative reduced cost violates (18) and can therefore be added to the set of independent sets in RCP $\Lambda^{\prime}$.
@@
$\begin{split}\sigma&=\Gamma_{I}-\sum\limits_{k}a^{k}_{I}\mu^{\prime}_{k}\\ &=\sum\limits_{k}a_{I}^{k}\Gamma_{c}^{k}-\sum\limits_{k}a^{k}_{I}\mu^{\prime}_{k}\\ &=\sum\limits_{k}a_{I}^{k}(\Gamma_{c}^{k}-\mu^{\prime}_{k})\end{split}$ (20)
@@
$\max\sigma=\sum\limits_{k}a_{I}^{k}(\mu^{\prime}_{k}-\Gamma_{c}^{k})$ (21)
By definition, $(a_{I}^{k})_{k\in K}$ is a vector denoting an independent set, and $\mu^{\prime}_{k}-\Gamma_{c}^{k}$ can be seen as numerical weights for every color $c$. This is therefore a Maximum Weighted Independent Set problem. For the rest of the paper, this problem will be our worker problem (WP). The MWIS problem can be defined as follows.
$\max\sigma=$ $\sum_{k}y_{k}(\mu^{\prime}_{k}-\Gamma_{c}^{k})$ (WP)
s.t. $y_{k}+y_{j}\leq 1$ $(\forall(k,j)\in E)$ (WP-1)
$y_{k}\in\{0,1\}$ $(\forall k\in K)$ (WP-2)
A solution of Problem (WP) defines an allocation $(a_{I},c)$ where $I\coloneqq\{k\in K|y_{k}=1\}$ and $a_{I}^{k}=y_{k}$. Note that there is one worker problem per color, and, as in our problem, colors from the same model are equivalent, there is actually one problem per vehicle model. Therefore, each allocation produced by a worker represents a single physical vehicle (color) along with all the tours assigned to it. Given the above problem definitions, we can define the algorithm to solve RCP as described in Algorithm 1.